# pFedGPA: Diffusion-based Generative Parameter Aggregation for Personalized Federated Learning

### v1

@@ Introduction
where $\theta^{*}$ represents the globally optimal parameters. The representative method for solving Eq. (1) is FedAvg ((McMahan et al. 2017b)). In each round, clients perform several epochs of SGD on their local loss functions and send the updated models to the server. The server then aggregates the models from all clients linearly by
$\bar{\theta}=\sum_{i=1}^{n}\frac{m_{i}}{N}\theta^{i}$ (2)
and broadcasts the averaged model $\bar{\theta}$ back to the clients.
@@
Forward process
Given training data $x_{0}$ from a target distribution $q(x_{0})$, the forward process gradually adds Gaussian noise to diffuse $x_{0}$ into $x_{1},x_{2},\dots,x_{T}$, where $x_{T}$ is approximately sampled from the standard Gaussian distribution $q(x_{T})\approx\mathcal{N}(x_{T};\mathbf{0},\mathbf{I})$. This process can be formulated as follows:
$q(x_{1:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1})$ (4)
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ (5)
where $\beta_{1},\dots,\beta_{T}\in(0,1)$ is a variance schedule. The diffusion model aims to approximate the noise added at each time step $t$ using a neural estimator $\epsilon_{\phi}(x_{t},t)$. The training objective $L_{ddpm}$ is given by the Eq. (6) below:
$L_{ddpm}=\mathbb{E}_{t\sim[1,T],x_{0}\sim q_{0},\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\|\epsilon_{t}-\epsilon_{\phi}(x_{t},t)\|_{2}^{2}],$ (6)
where $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ and $\bar{\alpha}_{t}={\textstyle\prod_{j=1}^{t}}(1-\beta_{j})$.
Reverse process
During the reverse process, we iteratively reconstruct the target data $x_{0}$ from the random noise $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. At each time step $t$, we sample $x_{t-1}$ from the estimated reverse Markov chain $p_{\phi}(x_{t}|x_{t-1})$ parameterized by the noise estimator $\epsilon_{\phi}(x_{t},t)$. That is,
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{(1-\overline{\alpha}_{t})}}\epsilon_{\phi}(x_{t},t))+\sigma_{t}z_{t}.$ (7)
@@
$\mathbb{E}_{q(x_{0})}\mathbb{E}_{q(x_{t}|x_{0})}[\|s_{\phi}(x_{t},t)-\nabla_{x_{t}}\log{q(x_{t}|x_{0})}\|_{2}^{2}],$ (8)
where $s_{\phi}(x_{t},t)=-\frac{\epsilon_{\phi}(x_{t},t)}{\sqrt{(1-\overline{\alpha}_{t})}}$. This allows us to use DDPMs to effectively estimate the scores of data distributions.
@@
In this section, we present our framework pFedGPA in detail. Given a collection of clients with their own specific data distributions in a FL system, our goal is to integrate the diverse parameter distributions captured by the diffusion model on the server, which reflects the underlying data distributions of the clients. Meanwhile, the diffusion model aims to generate new, personalized model parameters for each client by leveraging these learned parameter distributions. In the following subsections, we first illustrate the parameter integration process using the diffusion model. Next, we explain the personalized parameter generation process. Finally, we delve into the architecture design of the diffusion model. Further details about the training procedure are presented in Algorithm 1.
Figure 2: Illustration of the training process for a new client $k$ with and without global guidance. The blue arrows represent training starting from the initial parameters $\theta^{k}_{0}$ solely on local data $D_{k}$, gradually converging to the final optimized parameters $\theta^{k}_{*}$ within $S$ iterations. The orange arrows indicate training with global guidance $G$ alternated with local data training, which accelerates initialization and converges within $I\ll S$ iterations.
@@
In our framework, we deploy a diffusion model on the server, where both the input and output are the weights of the clients’ models. The core of this approach is to enable the server to learn the distribution of all clients’ model parameters using a powerful generative model, a distribution which resides on a low-dimensional manifold in the parameter space.
@@
where $\mathcal{L}_{i}(\phi,\theta^{i})=\mathbb{E}_{(x,y)\sim P_{i}}[\ell_{i}(F_{i}(H(\phi;\theta^{i});x),y)]$ and $\{\theta^{i}_{*}\}_{i=1}^{n}$ are the stationary points of the system $H$ under well-fitted state $\phi^{*}$ that represent the optimal model parameters of each client.
Moreover, in (Shamsian et al. 2021), the hypernetwork takes a trainable vector descriptor as the input to differentiate each client. However, This approach does not scale efficiently. When new clients join the network, it often requires training suitable embedding vectors for them from scratch. As the network scale continues to grow, it also becomes necessary to fine-tune all clients’ embedding vectors to achieve stronger representational capabilities while keeping consistency across clients. In diffusion models, if the input also includes client descriptors $e_{client}$, it can be implemented as a form of classifier-free guidance, utilizing a linear combination of conditional and unconditional score estimates to provide client-specific guidance:
@@
where $\omega$ controls the strength of the guidance. The advantage is no need to train a separate classifier. However, here we adopt a classifier guidance approach because, in the FL setting, each client inherently serves as a classifier model relative to the diffusion model. In turn, the diffusion model provides global guidance to local clients. That is:
@@
where $\log P_{y}(x|\theta)$ represents the training loss of the client model in probabilistic form (e.g., cross-entropy loss). Another benefit is that utilizing a unified model to estimate the overall score of model distributions, rather than conditionally fitting each client’s distribution, promotes information sharing across clients while maintaining the adaptability of personalized models. For a newly joined client, this method can accomplish initialization within a few iterations. During each iteration, the client sends fine-tuned model parameters to the server, where the diffusion model alternately provides global guidance to update them. This process is depicted in Fig.  2.
@@
For existing clients, we aim to generate parameters that align with their local data distributions while incorporating patterns learned from the diffusion model, ensuring the new parameters effectively fit the original data. Inspired by the inversion technique (Wu and De la Torre 2023) in image editing, we propose a parameter inversion method. We view the above difficulty as a zero-shot inversion problem within diffusion models. Given a parameter data, we can transform it into a latent code and then aggregate it cohesively with denoising sampling.
Detailed explanation is as follow. First, in the forward process of diffusion algorithm, Gaussian noises are added at each time step, diffusing $\theta_{0}$ to $\theta_{1},\dots,\theta_{T}$. Next, We concatenate all the added noises with the final $\theta_{T}$ to define a latent code of $\theta_{0}$. Consequently, the latent space has a much higher dimensionality than the original data. The formulations are as follows:
@@
where $\mathbf{\gamma}$ represents the implicit semantics of $\theta_{0}$. We can perfectly reconstruct $\theta_{0}$ from the latent code $\mathbf{\gamma}$:
@@
We use the latent diffusion model (Rombach et al. 2022) as the generative model and adopt the main architecture from p-diff (Wang et al. 2024). The latent diffusion model contains an autoencoder and a diffusion model. First, the autoencoder is trained through the reconstruction loss, producing a low-dimensional latent space. Then, the diffusion model operates within this latent space, learning to progressively denoise samples from noise to approximate the distribution of the latent representations. This approach allows the diffusion model to generate high-quality samples with reduced computational costs and memory usage compared to operating directly in the original high-dimensional space.
We flatten the model parameters into one-dimensional vectors, and both the U-Net blocks within the diffusion model and the encoder and decoder in the autoencoder use 1-D convolution layers instead of the traditional 2-D convolution layers. In addition, we introduced noise into the input and latent representations to improve the robustness and generalization of the generated results. Experiments in p-diff showed that noise augmentation plays a key role in producing models that are stable and achieve high performance. Unlike generating model parameters trained on IID data in p-diff, our diffusion model is designed to fit model parameters coming from different data distributions in FL. For smaller models, we consider generating the entire model, while for larger models, we generate specific layers due to hardware constraints.
@@
Datasets and Local Models
@@ Discussion
Communication and Privacy Concerns

### v2

@@ Introduction
where $\theta^{*}$ represents the globally optimal parameters. The representative method for solving Eq. (1) is FedAvg (McMahan et al. 2017b). In each round, clients perform several epochs of SGD on their local loss functions and send the updated models to the server. The server then aggregates the models from all clients linearly by $\bar{\theta}=\sum_{i=1}^{n}\frac{m_{i}}{N}\theta^{i}$ and broadcasts the averaged model $\bar{\theta}$ back to the clients.
@@
Forward process.
Given training data $z_{0}$ from a target distribution $q(z_{0})$, the forward process gradually adds Gaussian noise to diffuse $z_{0}$ into $z_{1},z_{2},\dots,z_{T}$, where $z_{T}$ is approximately sampled from the standard Gaussian distribution $q(z_{T})\approx\mathcal{N}(z_{T};\mathbf{0},\mathbf{I})$. This process can be formulated as follows:
$q(z_{1:T}|z_{0})=\prod_{t=1}^{T}q(z_{t}|z_{t-1})$ (3)
$q(z_{t}|z_{t-1})=\mathcal{N}(z_{t};\sqrt{1-\beta_{t}}z_{t-1},\beta_{t}\mathbf{I})$ (4)
where $\beta_{1},\dots,\beta_{T}\in(0,1)$ is a variance schedule. The diffusion model aims to approximate the noise added at each time step $t$ using a neural estimator $\epsilon_{\phi}(z_{t},t)$. The training objective $L_{ddpm}$ is given by the Eq. (5) below:
$L_{ddpm}=\mathbb{E}_{t\sim[1,T],z_{0}\sim q_{0},\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\|\epsilon_{t}-\epsilon_{\phi}(z_{t},t)\|_{2}^{2}],$ (5)
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$ and $\bar{\alpha}_{t}={\textstyle\prod_{j=1}^{t}}(1-\beta_{j})$.
Reverse process.
During the reverse process, we iteratively reconstruct the target data $z_{0}$ from the random noise $z_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. At each time step $t$, we sample $z_{t-1}$ from the estimated reverse Markov chain $p_{\phi}(z_{t}|z_{t-1})$ parameterized by the noise estimator $\epsilon_{\phi}(z_{t},t)$. That is,
$z_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(z_{t}-\frac{\beta_{t}}{\sqrt{(1-\overline{\alpha}_{t})}}\epsilon_{\phi}(z_{t},t))+\sigma_{t}z_{t}.$ (6)
@@
$\mathbb{E}_{q(z_{0})}\mathbb{E}_{q(z_{t}|z_{0})}[\|s_{\phi}(z_{t},t)-\nabla_{z_{t}}\log{q(z_{t}|z_{0})}\|_{2}^{2}],$ (7)
where $s_{\phi}(z_{t},t)=-\frac{\epsilon_{\phi}(z_{t},t)}{\sqrt{(1-\overline{\alpha}_{t})}}$. This allows us to use DDPMs to effectively estimate the scores of data distributions.
@@
In this section, we present our framework pFedGPA in detail. Given a collection of clients with their own specific data distributions in a FL system, our goal is to integrate the diverse parameter distributions captured by the diffusion model on the server, which reflects the underlying data distributions of the clients. Meanwhile, the diffusion model aims to generate new, personalized model parameters for each client by leveraging these learned parameter distributions. In the following subsections, we first illustrate the parameter aggregation using the diffusion model. Next, we explain the personalized parameter generation. Finally, we delve into the architecture design of the diffusion model. More details about the training procedure are presented in Algorithm 1.
Figure 2: Illustration of the training process for a new client $k$ with and without global guidance. The green arrows represent training starting from the initial parameters $\theta^{k}_{0}$ solely on local data $D_{k}$, gradually converging to the final optimized parameters $\theta^{k}_{*}$ within $S$ iterations. The purple arrows indicate training with global guidance $G$ alternated with local data training, which accelerates initialization and converges within $I\ll S$ iterations.
@@
In our framework, we deploy a diffusion model on the server, where both input and output are the weights of the clients’ models. The core of this approach is to enable the server to learn the distribution of all clients’ model parameters using a powerful generative model, a distribution which resides on a low-dimensional manifold in the parameter space.
@@
where $\mathcal{L}_{i}(\phi,\theta^{i})=\mathbb{E}_{(x,y)\sim P_{i}}[\ell_{i}(F_{i}(H(\phi;\theta^{i});x),y)]$ and $\{\theta^{i}_{*}\}_{i=1}^{n}$ are the stationary points of the system $H$ under well-fitted state $\phi^{*}$ that represent the optimal model parameters of each client. In each round, the server receives the client parameters and treats them as training data for the diffusion model, training it over a predefined number of epochs. By performing parameter inversion, which will be introduced in the next section, the server then generates personalized model parameters for each client.
Global Guidance.
To quickly initialize a newly joined client, we propose a global guidance approach to help the client model adapt its parameters to the local data in just a few iterations. In (Shamsian et al. 2021), a hypernetwork is used to take a trainable vector descriptor as input to differentiate each client. However, this approach has scalability limitations. Specifically, when new clients join the network, it often requires training new embedding vectors from scratch, which becomes increasingly inefficient as the network grows. Moreover, fine-tuning all clients’ embedding vectors to achieve consistent representation is necessary but complex. In diffusion models, if the input also includes client descriptors $e_{client}$, it can be implemented as a form of classifier-free guidance, utilizing a linear combination of conditional and unconditional score estimates to provide client-specific guidance:
@@
where $\omega$ controls the strength of the guidance. The advantage is that there is no need to train a separate classifier. However, in our approach, we adopt classifier-based guidance because, in the FL setting, each client inherently serves as a classifier relative to the diffusion model. In turn, the diffusion model offers global guidance to the local client models. That is:
@@
where $\log P_{y}(x|\theta)$ represents the training loss of the client model in probabilistic form (e.g., cross-entropy loss). This setup allows the diffusion model to share information across clients while preserving the adaptability of personalized models. Instead of fitting each client’s distribution individually, a unified model is used to estimate the overall score of model distributions, thereby facilitating efficient knowledge sharing.
For newly joined clients, this approach enables rapid initialization within just a few iterations. In each iteration, the client sends its fine-tuned model parameters to the server, where the diffusion model provides global guidance to update these parameters. This process is illustrated in Fig. 2.
@@
Generating personalized parameters remains challenging because a single diffusion model cannot directly control outputs for individual clients. In image generation, control is often achieved through conditioning on labels, but in our scenario, no such labels are available. To address this, we propose a novel parameter inversion method inspired by unsupervised inversion techniques (Wu and De la Torre 2023). In our approach, the uploaded model parameters serve as implicit ’labels’, which are decomposed into latent codes and injected into the diffusion model to generate stable, personalized parameters for each client.
Detailed explanation is as follows. First, in the forward process of the diffusion algorithm, Gaussian noises are added at each time step, diffusing $\theta_{0}$ to $\theta_{1},\dots,\theta_{T}$. Next, we concatenate all the added noises with the final $\theta_{T}$ to define a latent code of $\theta_{0}$. The formulations are as follows:
@@
where the latent code $\mathbf{\gamma}$ encodes the implicit semantics of $\theta_{0}$. We can perfectly reconstruct $\theta_{0}$ from $\mathbf{\gamma}$:
@@
We use the latent diffusion model (Rombach et al. 2022) as the generative model and adopt the main architecture from p-diff (Wang et al. 2024). The latent diffusion model consists of an autoencoder and a diffusion model. The autoencoder is first trained using reconstruction loss to produce a low-dimensional latent space. Then, the diffusion model operates within this latent space, learning to progressively denoise samples from noise, thus approximating the distribution of the latent representations. This approach enables the diffusion model to generate high-quality samples with reduced computational costs and memory usage compared to operating directly in the original high-dimensional space.
We flatten the model parameters into one-dimensional vectors and use 1-D convolution layers within both the encoder and decoder of the autoencoder, as well as the diffusion model, instead of the traditional 2-D convolution layers. Additionally, we introduce noise augmentation into the input and latent representations to improve the robustness and generalization of the generated results. Experiments in p-diff show that this augmentation is crucial for producing stable models with high performance. Unlike generating model parameters trained on IID data in p-diff, our diffusion model is designed to fit model parameters from different data distributions in FL. For smaller models, we generate the entire model, while for larger models, we generate specific layers due to hardware constraints. The complete details of the model are provided in the Appendix.
@@
Datasets and Local Models.
@@ Discussion
Communication and Privacy Concerns.
@@ Conclusion
Acknowledgments
This work is supported in part by the Natural Science Foundation of China (Grant 62371270) and Shenzhen Key Laboratory of Ubiquitous Data Enabling (No.ZDSYS20220527171406015).